Points A, B, C, and D are arranged in order on a line so that AB = 3BC and BC = 2CD. What fraction of AD is BD?
Answer by edjones(8007)

Let CD=1
Then BC=2 and AB=6
AD=9 and BD=3
So BD/AD=3/9=1/3
Ed